OpenStudy (mikey):

I'm so lost What is the simplified form of the seventh root of x to the fifth power times the seventh root of x to the fifth power? the square root of x x to the four ninth power x x times the seventh root of x cubed

OpenStudy (jiteshmeghwal9):

\[\Large(\sqrt[7]{x})^5*(\sqrt[7]{x})^5=(x^{\frac{5}{7}})^2\]

563blackghost (563blackghost):

Great explanation here :) http://brainly.com/question/1527386

563blackghost (563blackghost):

@mikey

OpenStudy (mikey):

sorry my computer restarted for updates